# Fine-grained and Explainable Factuality Evaluation for Multimodal Summarization

### v2

@@ Abstract
Fine-grained and Explainable Factuality Evaluation
for Multimodal Summarization
Yue Zhang1, Jingxuan Zuo2, Liqiang Jing1 1University of Texas at Dallas 2Shandong University
1 Introduction
@@
Sentence summarization, a pivotal task in natural language processing, focuses on creating concise summaries from longer sentences. This area has gained significant attention due to its utility in summarizing events and products, as indicated by various studies. Traditionally, this summarization relied solely on the source sentence. However, with the growing presence of multimedia content combining text and images, recent research has evolved towards multimodal summarization (MMS) Song et al. (2022); Jing et al. (2023); Lin et al. (2023). This approach integrates visual cues with textual information to enhance the summarization process, making it easier for readers to quickly capture the essence of the information.
@@
In the realm of MMS, notable advancements have been made. Li et al. employed sequence-to-sequence models for better semantic understanding and text generation. They further enhanced MMS by introducing a multimodal selective gate network, which helped in pinpointing the most relevant parts of a sentence based on its corresponding image. More recently, Song et al. developed a method for generating efficient product summaries using generative pre-trained language models like BART Lewis et al. (2020). This method begins by transforming product images into attributes, which are then used by the BART model for generating succinct summaries. This shift towards incorporating visual elements marks a significant development in the field of sentence summarization. A significant challenge faced by these models is the issue of hallucination Wan and Bansal (2022a). This occurs when the model produces content that is neither present nor implied in the original input text and image.
@@
Recent progress has been made in creating metrics that align closely with human assessments of factual accuracy in summaries Tang et al. (2022); Zhu et al. (2021). These metrics are designed to evaluate the level of factual consistency between the original document and its generated summary. However, there is only one related work that employs factual accuracy in the multimodal summarization task. Wan and Bansal proposed CLIPBERTSCORE, a simple weighted combination of CLIPScore Hessel et al. (2021) and BERTScore Zhang et al. (2020a) to leverage the robustness and strong factuality detection performance between image-summary and document summary, respectively.
@@
However, BERTScore pays more attention to more grammatically correct sentences rather than factual sentences Hanna and Bojar (2021). Therefore, the factuality of the text input cannot be measured. In addition, due to CLIPScore’s limitations in accurately counting objects Radford et al. (2021) or conducting compositional reasoning Ma et al. (2023), the CLIPScore often proves to be unreliable and can yield inaccurate results. Furthermore, the existing evaluation metric only generates an overall score, which is coarse-grained and less explainable.
@@
2 Related Work
Multimodal Summarization Recently, multimodal summarization, which integrates text with images, has emerged as a significant advancement over traditional text summarization Li et al. (2020b, a); Zhang et al. (2021); Li et al. (2020c); Palaskar et al. (2019). Early works focused on combining Convolutional Neural Network-based visual models with Recurrent Neural Network-based textual models to improve the multimodal summarization Zhu et al. (2018); Chen and Zhuge (2018); Li et al. (2018); Zhu et al. (2020); Zhang et al. (2022). Recent research in the field has shifted towards leveraging pre-trained models for multimodal summarization Jiang et al. (2023); Lin et al. (2023); Jing et al. (2023); Song et al. (2022). Despite the superior performance of the existing models, these models always suffer from the hallucination where the model generates contents that are not present or entailed by inputs Wan and Bansal (2022a).
Faithfulness and Factuality Metrics Evaluating the factuality of generated summaries is crucial for their reliability and accuracy. Factuality metrics fall into two main categories: entailment-based metrics and question-answering-based (QGQA) metrics. Entailment-based metrics Kryscinski et al. (2020); Goyal and Durrett (2021) assess whether summaries reflect the source document accurately. QGQA approaches Durmus et al. (2020); Wang et al. (2020); Scialom et al. (2021); Fabbri et al. (2022), evaluate the factuality of the summary by generating and answering input-related questions. Additionally, some other efforts, such as counterfactual estimation Xie et al. (2021) and embedding-based metrics Zhang et al. (2020b) have been introduced to this field. Different from them, Wan and Bansal is the first one to consider the textual and visual modalities for multimodal summarization factuality evaluation. Although they have achieved compelling success, these methods do not take into account fine-grained elements and are hard to explain. To address this issue, we introduced a framework involving entity-focused questions to evaluate the model’s fine-grained factuality, which allows for a more detailed assessment of how accurately the model reflects specific entities and relations within the multimodal input.
3 Frameworks
3.1 Problem Definition
@@ 3.2.1 Reference-based
To get a fine-grained evaluation in terms of factuality, we first generate a question set $\mathcal{Q}$ based on the input document, which can generate comprehensive evaluation questions. All the questions are based on atomic information Min et al. (2023). We give some examples in Appendix D. The prompt for question generation is detailed in Appendix C. This process is implemented by a Large Language Model (LLM), formalized as $LLM(Y,D)\rightarrow\mathcal{Q}$.
@@ 3.2.2 Reference-free
To verify the consistency between the multimodal input and summary, we compute the consistency for each modality. Specifically, we use a QA model to answer questions $\mathcal{Q}$ based on the document $D$ and a VQA model to answer questions $\mathcal{Q}$ based on the image $I$. The QA model also is GPT-4, and the VQA model is implemented by BLIP-2 Li et al. (2023a). We denote the set of answers as $\mathcal{A}_{t}$ and $\mathcal{A}_{i}$ corresponding to the document $D$ and the image $I$.
@@
where $\mathbb{I}$ denotes the indicator function, $H(x)$ represents a condition where $x$ is equal to ’Yes’, $\lor$ symbolizes the logical OR operation, and $|\mathcal{Q}|$ is the total number of questions in $\mathcal{Q}$.
In this formulation, we consider that if either the $a_{t}$ or the $a_{i}$ contains ’Yes’, it indicates factuality, signifying that the model has not fabricated information. Otherwise, it indicates a hallucination.
4 Experiments
4.1 Experimental Setting
We employed the BART-MMSS Lin et al. (2023) model and MMSS dataset Li et al. (2018) to evaluate the effectiveness of our proposed method. We employed GPT-4 to generate questions and as the QA model. We used BLIP-2 Li et al. (2023b) as the VQA model.
4.2 Human Evaluation
To verify the effectiveness of the reference-based evaluation framework, we randomly selected 200 samples from the MMSS dataset. In addition, we randomly annotated 200 samples generated by another model V2P Song et al. (2022) on another dataset CEPSUM Li et al. (2020a). More annotation details are detailed in Appendix A.
4.3 Result Analysis
Evaluating QA Model
Evaluating VQA Model
Comparison
We selected document-based metrics: 1) ROUGE-1 Lin (2004), 2) ROUGE-L Lin (2004), 3) BLEU Papineni et al. (2002), and 4) BertScore Zhang et al. (2020a); image-based metric: CLIPScore Hessel et al. (2021); and combined metric: comCLIPBERTScore Wan and Bansal (2022b) as baseline metrics. To the best of our knowledge, CLIPBERTScore has been identified as the best metric for evaluating faithfulness in previous assessments of multi-modal summarization. Besides, we also conduct different ablation methods derived from our frameworks based on image input and document input, respectively.
We computed Pearson correlation coefficients between the human evaluation results and all atomic metrics (all baselines and our proposed metrics) in terms of factuality. From Table 1 and 2, we have two observations: (1) Compared with document-based metrics, image-based metrics always perform worse. The potential reason may be that the summary has more overlapped content with the input text rather than the input image. (2) Our frameworks achieved the best performance compared to other baselines, which demonstrates the superiority of our method. In addition, we also use the ChatGPT for evaluation on MMSS and achieve a 0.48/0.47 Pearson correlation coefficient for reference-free/based metric which surpasses all the baselines and demonstrates our framework has the generalization ability.
@@ 4.3 Result Analysis
5 Conclusion
@@
Limitations
Currently, our work relies on Large Language Models (i.e., GPT-4) for evaluation, which is expensive when you evaluate massive datasets. In addition, we used the changing LLMs and the black-box model, which may cause difficulties in reproductivity. The LLMs used in the evaluation metrics may introduce propagated bias.
Appendix A Details of Human Evaluation
For the dataset MMSS, each sample has a maximum of 12 QA pairs, and a minimum of 4 QA pairs, with an average of 8.34 QA pairs per text, totaling 1668 QA pairs. On average, the length of the source sentence and the target summary are 22 and 8, respectively. Similarly, for the reference-free evaluation framework, we also utilized these 200 samples. For this framework, the number of generated questions is 3. For the dataset CEPSUM, each sample has a maximum of 15 QA pairs, and a minimum of 5 QA pairs, with an average of 10.2 QA pairs per text, totaling 2040 QA pairs. The average token lengths of the input document and output summary are 316 and 79, respectively. Similarly, for the reference-free evaluation framework, we also utilized these 200 samples. For this framework, the number of generated questions is 6.
Appendix B More Analysis
For the reference-based framework, we found the ground truth replaces words or rephrases input documents, which makes the generated questions more difficult compared to the reference-free framework. This may cause the performance of a reference-free framework to be better than a reference-based framework. Compared to text, images show a low correlation with humans, which has also been found in another dataset Wan and Bansal (2022a). Therefore, we argue that our method can be applied to more datasets. In addition, results on another model V2P, and another dataset CEPSUM show our method is reliable and can be transferred to other datasets.
CLIPBERTScore is a simple weighted combination of CLIPSCORE and BERTSCore. However, it’s important to note that while BERTScore tends to focus more on the grammatical correctness of sentences, it may not necessarily prioritize factual accuracy Hanna and Bojar (2021). Furthermore, CLIPScore has its limitations, notably in accurately modeling entities Radford et al. (2021) and in conducting compositional reasoning Ma et al. (2023). These limitations can lead to reliability issues, often resulting in inaccurate outcomes. As a result, the effectiveness of CLIPBERTScore in delivering accurate factuality scores is potentially compromised. In contrast, our proposed metric offers a more nuanced approach. It is not only fine-grained, allowing for a detailed analysis, but also interpretable, providing clear insights into its scoring process.
Appendix C Prompt
We illustrate our prompt for question generation in Table 3.
Table 3: Prompt for question generation and question answering.
Question Generation for the Task of With Ground Truth: User: Now, you will receive a news article. Please generate questions and provide answers based on the information in the news article. Your questions should be proportional to the amount of information and the number of named entities in the article. The question should be a Yes-No Question.The answers should be "yes," "no," or "not provided," and the answer is based on news article. In other words, you do not need to ask open-ended questions. All questions must have answers, either "yes," "no," or "not provided." Please intentionally ask some incorrect questions to verify the accuracy of the answers. For example, you can intentionally create incorrect answers using names or places that have not appeared before to obtain QA pairs with answers as "NO.". Try to maintain a roughly equal ratio of "yes," "no," and "not provided" answers and cover as much content as possible. The format for providing answers should be ["Question": question, "Answer": answer, "Question": question, "Answer": answer, "Question": question, "Answer": answer, "Question": question, "Answer": answer].The news article is:"u.s. , albanian , croatian and macedonian soldiers began a monthlong joint military exercise tuesday , the u.s. embassy said tuesday ." GPT: [ "Question": "Did the joint military exercise begin on Tuesday?", "Answer": "Yes" , "Question": "Is the joint military exercise only between the U.S. and Albania?", "Answer": "No" , "Question": "Is Croatian military involved in the joint exercise?", "Answer": "Yes" , "Question": "Did the French military participate in the joint exercise?", "Answer": "No" , "Question": "Was the announcement made by the U.S. Embassy?", "Answer": "Yes" , "Question": "Did the joint military exercise last for a year?", "Answer": "No" , "Question": "Did Macedonian soldiers participate in the joint exercise?", "Answer": "Yes" , "Question": "Did the news summary provide information about the purpose of the exercise?", "Answer": "No" , "Question": "Did the exercise commence on Wednesday?", "Answer": "No" , "Question": "Is the joint exercise expected to last for a month?", "Answer": "Yes" ] Question Generation for the Task of Without Ground Truth: User: I will now provide you with a news summary. Please generate three yes/no questions regarding the entities and relationships within this sentence. The answers to these three questions should all be true. Please present the questions in the format of ["question1", "question2", "question3"].The summary is:u.s. albanian croatian and macedonian soldiers joint military exercise.s. GPT: [ "Are U.S. soldiers participating in a joint military exercise with Albanian soldiers?", "Are U.S. soldiers participating in a joint military exercise with Croatian soldiers?", "Are U.S. soldiers participating in a joint military exercise with Macedonian soldiers?" ] Question Answering: User:I will provide you with a news segment and a question; please provide the answers to the questions in the form of 0 or 1, where 0 represents no and 1 represents yes. news:u.s. , albanian , croatian and macedonian soldiers began a monthlong joint military exercise Tuesday, the u.s. embassy said Tuesday. question: Are U.S. soldiers participating in a joint military exercise with Albanian soldiers? GPT:1
Appendix D Evaluation Example
We presented reference-free and reference-based evaluation metrics in Figure 3 and Figure 4.
Figure 3: The reference-based evaluation example.
Figure 4: The reference-free evaluation example. QA denotes the answer based on textual modality and VQA denotes the answer based on visual modality. The final answer is the combination of QA and VQA. Score means the score of the proposed metric

### v3

@@ Abstract
Introduction
@@
Sentence summarization, a pivotal task in natural language processing, focuses on creating concise summaries from longer sentences. This area has gained significant attention due to its utility in summarizing events and products, as indicated by various studies. Traditionally, this summarization relied solely on the source sentence. However, with the growing presence of multimedia content combining text and images, recent research has evolved towards multimodal summarization (MMS) (Song et al. 2022; Jing et al. 2023; Lin et al. 2023). This approach integrates visual cues with textual information to enhance the summarization process, making it easier for readers to quickly capture the essence of the information.
@@
In the realm of MMS, notable advancements have been made. Li et al. employed sequence-to-sequence models for better semantic understanding and text generation. They further enhanced MMS by introducing a multimodal selective gate network, which helped in pinpointing the most relevant parts of a sentence based on its corresponding image. More recently, Song et al. developed a method for generating efficient product summaries using generative pre-trained language models like BART (Lewis et al. 2020). This method begins by transforming product images into attributes, which are then used by the BART model for generating succinct summaries. This shift towards incorporating visual elements marks a significant development in the field of sentence summarization. A significant challenge faced by these models is the issue of hallucination (Wan and Bansal 2022a). This occurs when the model produces content that is neither present nor implied in the original input text and image.
@@
Recent progress has been made in creating metrics that align closely with human assessments of factual accuracy in summaries (Tang et al. 2022; Zhu et al. 2021). These metrics are designed to evaluate the level of factual consistency between the original document and its generated summary. However, there is only one related work that employs factual accuracy in the multimodal summarization task. Wan and Bansal proposed CLIPBERTSCORE, a simple weighted combination of CLIPScore (Hessel et al. 2021) and BERTScore (Zhang et al. 2020a) to leverage the robustness and strong factuality detection performance between image-summary and document summary, respectively.
@@
However, BERTScore pays more attention to more grammatically correct sentences rather than factual sentences (Hanna and Bojar 2021). Therefore, the factuality of the text input cannot be measured. In addition, due to CLIPScore’s limitations in accurately counting objects (Radford et al. 2021) or conducting compositional reasoning (Ma et al. 2023), the CLIPScore often proves to be unreliable and can yield inaccurate results. Furthermore, the existing evaluation metric only generates an overall score, which is coarse-grained and less explainable.
@@
Frameworks
Problem Definition
@@
To get a fine-grained evaluation in terms of factuality, we first generate a question set $\mathcal{Q}$ based on the input document, which can generate comprehensive evaluation questions. All the questions are based on atomic information (Min et al. 2023). The prompt for question generation is detailed in Supplementary Material. This process is implemented by a Large Language Model (LLM), formalized as $LLM(Y,D)\rightarrow\mathcal{Q}$.
@@
To verify the consistency between the multimodal input and summary, we compute the consistency for each modality. Specifically, we use a QA model to answer questions $\mathcal{Q}$ based on the document $D$ and a VQA model to answer questions $\mathcal{Q}$ based on the image $I$. The QA model also is GPT-4, and the VQA model is implemented by BLIP-2 (Li et al. 2023a). We denote the set of answers as $\mathcal{A}_{t}$ and $\mathcal{A}_{i}$ corresponding to the document $D$ and the image $I$.
@@
where $\mathbb{I}$ denotes the indicator function, $H(x)$ represents a condition where $x$ is equal to ‘Yes’, $\lor$ symbolizes the logical OR operation, and $|\mathcal{Q}|$ is the total number of questions in $\mathcal{Q}$.
In this formulation, we consider that if either the $a_{t}$ or the $a_{i}$ contains ‘Yes’, it indicates factuality, signifying that the model has not fabricated information. Otherwise, it indicates a hallucination.
Experiments
@@
Experimental Setting
We employed the BART-MMSS (Lin et al. 2023) model and MMSS dataset (Li et al. 2018) to evaluate the effectiveness of our proposed method. We employed GPT-4 to generate questions and as the QA model. We used BLIP-2 (Li et al. 2023b) as the VQA model.
Human Evaluation
To verify the effectiveness of the reference-based evaluation framework, we randomly selected 200 samples from the MMSS dataset. In addition, we randomly annotated 200 samples generated by another model V2P (Song et al. 2022) on another dataset CEPSUM (Li et al. 2020a). For the dataset MMSS, each sample has a maximum of 12 QA pairs, and a minimum of 4 QA pairs, with an average of 8.34 QA pairs per text, totaling 1668 QA pairs. On average, the length of the source sentence and the target summary are 22 and 8, respectively. Similarly, for the reference-free evaluation framework, we also utilized these 200 samples. For this framework, the number of generated questions is 3. For the dataset CEPSUM, each sample has a maximum of 15 QA pairs, and a minimum of 5 QA pairs, with an average of 10.2 QA pairs per text, totaling 2040 QA pairs. The average token lengths of the input document and output summary are 316 and 79, respectively. Similarly, for the reference-free evaluation framework, we also utilized these 200 samples. For this framework, the number of generated questions is 6.
Result Analysis
Evaluating QA Model
Evaluating VQA Model
(a) The reference-based evaluation example
(b) The reference-free evaluation example
Figure 3: Examples of evaluation methods. QA denotes the answer based on textual modality and VQA denotes the answer based on visual modality. The final answer is the combination of QA and VQA. Score means the score of the proposed metric.
Comparison
We selected document-based metrics: 1) ROUGE-1 (Lin 2004), 2) ROUGE-L (Lin 2004), 3) BLEU (Papineni et al. 2002), and 4) BertScore (Zhang et al. 2020a); image-based metric: CLIPScore (Hessel et al. 2021); and combined metric: comCLIPBERTScore (Wan and Bansal 2022b) as baseline metrics. To the best of our knowledge, CLIPBERTScore has been identified as the best metric for evaluating faithfulness in previous assessments of multi-modal summarization. Besides, we also conduct different ablation methods derived from our frameworks based on image input and document input, respectively.
We computed Pearson correlation coefficients between the human evaluation results and all atomic metrics (all baselines and our proposed metrics) in terms of factuality. From Table 1 and 2, we have two observations: (1) Compared with document-based metrics, image-based metrics always perform worse. The potential reason may be that the summary has more overlapped content with the input text rather than the input image. (2) Our frameworks achieved the best performance compared to other baselines, which demonstrates the superiority of our method. In addition, we also use the ChatGPT for evaluation on MMSS and achieve a 0.48/0.47 Pearson correlation coefficient for reference-free/based metric which surpasses all the baselines and shows our framework has the generalization ability. We presented reference-free and reference-based evaluation metrics in Figure 3(a) and 3(b).
More Analysis
For the reference-based framework, we found the ground truth replaces words or rephrases input documents, which makes the generated questions more difficult compared to the reference-free framework. This may cause the performance of a reference-free framework to be better than a reference-based framework. Compared to text, images show a low correlation with humans, which has also been found in another dataset (Wan and Bansal 2022a). Therefore, we argue that our method can be applied to more datasets. In addition, results on another model V2P, and another dataset CEPSUM show our method is reliable and can be transferred to other datasets.
CLIPBERTScore is a simple weighted combination of CLIPSCORE and BERTSCore. However, it’s important to note that while BERTScore tends to focus more on the grammatical correctness of sentences, it may not necessarily prioritize factual accuracy (Hanna and Bojar 2021). Furthermore, CLIPScore has its limitations, notably in accurately modeling entities (Radford et al. 2021) and in conducting compositional reasoning (Ma et al. 2023). These limitations can lead to reliability issues, often resulting in inaccurate outcomes. As a result, the effectiveness of CLIPBERTScore in delivering accurate factuality scores is potentially compromised. In contrast, our proposed metric offers a more nuanced approach. It is not only fine-grained, allowing for a detailed analysis, but also interpretable, providing clear insights.
Related Work
Multimodal Summarization Recently, multimodal summarization, which integrates text with images, has emerged as a significant advancement over traditional text summarization (Li et al. 2020b, a; Zhang et al. 2021; Li et al. 2020c; Palaskar et al. 2019). Early works focused on combining Convolutional Neural Network-based visual models with Recurrent Neural Network-based textual models to improve the multimodal summarization (Zhu et al. 2018; Chen and Zhuge 2018; Li et al. 2018; Zhu et al. 2020; Zhang, Zhang, and Pan 2022). Recent research in the field has shifted towards leveraging pre-trained models for multimodal summarization (Jiang et al. 2023; Lin et al. 2023; Jing et al. 2023; Song et al. 2022). Despite the superior performance of the existing models, these models always suffer from the hallucination where the model generates contents that are not present or entailed by inputs (Wan and Bansal 2022a).
Faithfulness and Factuality Metrics Evaluating the factuality of generated summaries is crucial for their reliability and accuracy. Factuality metrics fall into two main categories: entailment-based metrics and question-answering-based (QGQA) metrics. Entailment-based metrics (Kryscinski et al. 2020; Goyal and Durrett 2021) assess whether summaries reflect the source document accurately. QGQA approaches (Durmus, He, and Diab 2020; Wang, Cho, and Lewis 2020; Scialom et al. 2021; Fabbri et al. 2022), evaluate the factuality of the summary by generating and answering input-related questions. Additionally, some other efforts, such as counterfactual estimation (Xie et al. 2021) and embedding-based metrics (Zhang et al. 2020b) have been introduced to this field. Different from them, Wan and Bansal is the first one to consider the textual and visual modalities for multimodal summarization factuality evaluation. Although they have achieved compelling success, these methods do not take into account fine-grained elements and are hard to explain. To address this issue, we introduced a framework involving entity-focused questions to evaluate the model’s fine-grained factuality, which allows for a more detailed assessment of how accurately the model reflects specific entities and relations within the multimodal input.
Conclusion